OpenStudy (anonymous):

Which improper fraction is equivalent to 1 2/9

OpenStudy (anonymous):

12/9 11/9 19/9 17/9

OpenStudy (anonymous):

please help

jimthompson5910 (jim_thompson5910):

1 2/9 = 1 + 2/9 1 2/9 = 1(9/9) + 2/9 1 2/9 = 9/9 + 2/9 I'll let you finish

OpenStudy (anonymous):

?

jimthompson5910 (jim_thompson5910):

what do you get when you add 9/9 to 2/9

OpenStudy (anonymous):

11/9!!!!!!!!!!!!!!!!!!!!!!!!!

jimthompson5910 (jim_thompson5910):

you got it